6a.A rectangular airstrip measures 34.50 m by 230 m, with the width measured more accurately than the length. Find the area (in m2), taking into account significant figures.


1
Expert's answer
2022-06-06T08:54:48-0400
"A=a*b=34.50*230=7900\:\rm m^2"